OpenStudy (anonymous):

Differentiate each of the following functions with respect to x. (i) \[e^{x^3}\] (ii)\[{-1\over2}x^{(1/2)^2}\]

OpenStudy (anonymous):

Sorry, the (II) has no -1/2 infront

OpenStudy (anonymous):

\[\left( e ^{x ^{3}} \right)\prime ===> 3x ^{2}e ^{x ^{3}}\]

OpenStudy (anonymous):

Yes, but why?

OpenStudy (anonymous):

e^(x^3) ->e^(x^3) * (x^3)' = 3x^2 * e^(x^3)

OpenStudy (anonymous):

the derivative of e^x is itself times the derivative of x in this case x is x^3

OpenStudy (anonymous):

and derivative of X^3 is 3*x^2

OpenStudy (anonymous):

When you have, e^(something), the derivative is e^(something) multiplied by the derivative of something

OpenStudy (anonymous):

Ok, I understand. How about the second one?

OpenStudy (anonymous):

\[x^{(-1/2)^2}\]

OpenStudy (anonymous):

is that x^((-1/2)^2)

OpenStudy (anonymous):

tell me if im wrong but is that just -1/2 x^1/4 to simplify? then you just find that derivative

OpenStudy (anonymous):

sorry \[e^{(-1/2)x^2}\]

OpenStudy (anonymous):

e^n derivative is e^n * n' where n = (-1/2)*x^2 n'= -x

OpenStudy (anonymous):

\[(e ^{(-1/2)x ^{2}})\prime ===> e ^{(-1/2)x ^{2}}(-x)\]

OpenStudy (anonymous):

Oh. Ok

OpenStudy (anonymous):

derivative of e^(_x)=-e^(-x) ans=(-1/2)e^(-1/2)x^2{-x}=(1/2)e^(-1/2)x^2